a baseball is thrown to a distance of 890 m at an angle of 30 degree above horizontal. what is the initial speed of the baseball


1
Expert's answer
2022-01-30T13:51:34-0500

Given:

"R=890\;\rm m"

"\theta=30^\circ"


The range of projectile is given by

"R=\frac{v_0^2\sin 2\theta}{g}"

Hence, the initial velocity

"v_0=\sqrt{\frac{gR}{\sin 2\theta}}=\sqrt{\frac{9.8*890}{\sin 60^\circ}}=100\:\rm m/s"